OpenStudy (18jonea):

Explain in your own words how to find an inverse of a function. How can you use a graph to check to see if your two functions are in fact inverses? @3mar

OpenStudy (18jonea):

@ParthKohli

OpenStudy (3mar):

1. solve for x 2. swap x and y let's take ex

OpenStudy (3mar):

y+x+3=0 separate x in one side x=-3-y swap x and y y=-3-x Done!

OpenStudy (3mar):

@18jonea

OpenStudy (18jonea):

ok now how would you use a graph

OpenStudy (3mar):

That is easy Take that site. It is cool and interesting! https://www.desmos.com/calculator/zrnehzlbro

OpenStudy (3mar):

Thank you for the medal!

OpenStudy (18jonea):

ok but i have to explain in words how to do that

OpenStudy (3mar):

This is your mission should you choose to accept it ;) I tell the maths and you do the rest!

OpenStudy (18jonea):

i am not sure how to explain it though @3mar

OpenStudy (3mar):

So?

OpenStudy (18jonea):

like i know if they are opposites they are inverse of each other, but how would i use a graph?

OpenStudy (3mar):

http://www.purplemath.com/modules/invrsfcn3.htm very helpful link I think

OpenStudy (18jonea):

i am still not understanding

OpenStudy (3mar):

Let's express what we are doing!

OpenStudy (3mar):

If we have a function of x, let it is f(x). we should firstly check for inverse availability using horizontal line test. If it passes, the inverse is also a function! Then separate all x terms in one side and the other terms in the other side. i.e: solve for x Then replace every x with y and every y with x, so you got the function g(x) which is an inverse of the original function f(x) and you can illustrate by an example!

OpenStudy (18jonea):

ok so if they have reversed coordinantes, or if it passes the horizontal line test it is inverse?

OpenStudy (3mar):

If it passes horizontal line test, its inverse is also a function. If it does not pass the horizontal line test, the inverse is not a function!

OpenStudy (3mar):

Another expression of the inverse of a function is the image of this function when reflected across the line y=x. ...i.e graphically

OpenStudy (18jonea):

ok thank you! sorry for not underswtanding

OpenStudy (3mar):

Not at all! You are welcome!

OpenStudy (3mar):

Thank you for learning!!

OpenStudy (3mar):

Any more questions?

OpenStudy (18jonea):

not right now

OpenStudy (3mar):

Any Help... Anytime.